OpenStudy (anonymous):

Find the product of (3x + 7y)2. 9x2 - 42xy + 49y2 9x2 - 49y2 9x2 + 49y2 9x2 + 42xy + 49y2

OpenStudy (texaschic101):

(3x + 7y)^2 turns into (3x + 7y)(3x + 7y) Do you know how to do the FOIL method ?

OpenStudy (anonymous):

answer is D

OpenStudy (texaschic101):

FOIL -- first,outside,inside,last (3x + 7y)(3x + 7y) First -- multiply 1st terms in each set....3x * 3x = 9x^2 Outside-- multiply outside terms in each set...3x * 7y = 21xy Inside -- multiply inside terms in each set...7y * 3x = 21xy Last -- multiply last terms in each set....7y * 7y = 49y^2 now write what you got.. 9x^2 + 21xy + 21xy + 49y^2 combine like terms 9x^2 + 42xy + 49y^2

OpenStudy (texaschic101):

yep...you are correct....answer D

OpenStudy (anonymous):

thank can u help me on other questions

OpenStudy (texaschic101):

I can try...can't promise anything

OpenStudy (anonymous):

find the product of (x - 7)2. x2 - 14x + 49 x2 + 14x + 49 x2 - 49 x2 + 49

OpenStudy (texaschic101):

do it the same way...FOIL (x - 7)^2 is (x - 7)(x - 7) FOIL gives us : x^2 - 7x - 7x + 49 combine like terms x^2 - 14x + 49

OpenStudy (texaschic101):

anything else ?

OpenStudy (texaschic101):

I am sorry but I have to go now....getting kinda late.

OpenStudy (anonymous):

wait whats the answer

OpenStudy (texaschic101):

x^2 - 14x + 49

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can u help me on one more

OpenStudy (texaschic101):

no problem :)

OpenStudy (anonymous):

Find the product of (x + 3)2. x2 + 6x + 9 x2 - 6x + 9 x2 - 9 x2 + 9

OpenStudy (texaschic101):

yes...I guess I can help with 1 more

OpenStudy (texaschic101):

(x + 3)^2 is the same as (x + 3)(x + 3) using FOIL, we get : x^2 + 3x + 3x + 9 combine like terms x^2 + 6x + 9 <== answer

OpenStudy (anonymous):

thankuuuu

OpenStudy (anonymous):

Find the product of (3x - 7y)2. 9x2 - 42xy + 49y2 9x2 + 42xy + 49y2 9x2 - 49y2 9x2 + 49y2

OpenStudy (texaschic101):

sure thing :) Last one.. (3x - 7y)^2 (3x - 7y)(3x - 7y) FOIL gives us : 9x^2 - 21xy - 21xy + 49y^2 combine like terms 9x^2 - 42xy + 49y^2 <== answer

OpenStudy (texaschic101):

good-night now :)

OpenStudy (anonymous):

thanks wait

OpenStudy (anonymous):

Find the product of (x + 3)(x - 3). x2 - 6x + 9 x2 + 6x + 9 x2 + 9 x2 - 9

OpenStudy (texaschic101):

x^2 - 3x + 3x - 9 combine like terms x^2 - 9 <== answer

OpenStudy (anonymous):

okay i have one last one and im done

OpenStudy (anonymous):

Find the product of (3x + 7y)(3x - 7y). 9x2 - 42xy + 49y2 9x2 + 42xy + 49y2 9x2 - 49y2 9x2 + 49y2

OpenStudy (anonymous):

pleaseeee

OpenStudy (texaschic101):

9x^2 - 21xy + 21xy - 49y^2 combine like terms 9x^2 - 49y^2 <== your answer

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

no problem :)

OpenStudy (anonymous):

goodnight to u if u r awake let me no pls i need help